# DreamPBR: Text-driven Generation of High-resolution SVBRDF with Multi-modal Guidance

### v1

@@ Abstract.
Prior material creation methods had limitations in producing diverse results mainly because reconstruction-based methods relied on real-world measurements and generation-based methods were trained on relatively small material datasets. To address these challenges, we propose DreamPBR, a novel diffusion-based generative framework designed to create spatially-varying appearance properties guided by text and multi-modal controls, providing high controllability and diversity in material generation. Key to achieving diverse and high-quality PBR material generation lies in integrating the capabilities of recent large-scale vision-language models trained on billions of text-image pairs, along with material priors derived from hundreds of PBR material samples. We utilize a novel material Latent Diffusion Model (LDM) to establish the mapping between albedo maps and the corresponding latent space. The latent representation is then decoded into full SVBRDF parameter maps using a rendering-aware PBR decoder. Our method supports tileable generation through convolution with circular padding. Furthermore, we introduce a multi-modal guidance module, which includes pixel-aligned guidance, style image guidance, and 3D shape guidance, to enhance the control capabilities of the material LDM. We demonstrate the effectiveness of DreamPBR in material creation, showcasing its versatility and user-friendliness on a wide range of controllable generation and editing applications.
@@
Figure 1. DreamPBR, an innovative material generation framework, enables personalized creation with multi-modal controls. We presents various controls such as text descriptions (4, 5, 6, 7), binary images (2, 9, 10), RGB images (3, 11, 12, 13), segmented geometry (8) and thier combination (1) in this figure. The high-quality and tileable textures from DreamPBR show high applicability in different objects.
@@ 1. Introduction
Recent approaches have explored material generation (Guo et al., 2020; Zhou et al., 2022) using Generative Adversarial Networks (GANs) (Goodfellow et al., 2014). However, these methods are typically trained on hundreds to thousands of materials, which pales in comparison to the billions of images used in large-scale Language-Image generative models. The dataset capacity restricted their generating diversity. Furthermore, GAN-based methods also had training challenges including unstable training, mode collapse, and scalability issues with large datasets. On the other hand, diffusion models (Ho et al., 2020; Rombach et al., 2022) have shown significant advancements, exhibiting advantages in scalability and diversity. Recent advances (Poole et al., 2022; Wang et al., 2023) leverage 2D diffusion models as prior to generate 3D content. However, these methods mainly focus on implicit representation or textured mesh, lacking the capability to disentangle physically-based material and illumination.
@@
The key idea of our method is to integrate pretrained 2D text-to-image diffusion models (Rombach et al., 2022) with material priors to generate high-fidelity and diverse materials. While 2D text-to-image Latent Diffusion Models (LDM) excel in generating natural images, they had challenges to produce spatially-varying physically-based material maps due to the large domain gaps between natural images and materials. Consequently, adapting pretrained 2D diffusion models into material domain, while preserving both quality and diversity, is a non-trivial research task. We introduce a novel material LDM which are learned by a two-stage strategy to address this challenge. In the initial stage, we observed albedo map is specialized RGB image and stores spatially-varying surface reflectance by RGB pixel values. We transfer the pretrained LDM from text-to-image domain to text-to-albedo domain using fine-tuning, which can be regarded as the distillation from a large source domain (natural images) to a relatively small target domain (albedo texture maps) by leveraging the target domain priors. In the subsequent stage, we leverage a PBR decoder to reconstruct SVBRDFs from the latent space of albedo maps learned in the former stage. The reasons that we employ a decoder-only architecture for SVBRDFs generation are: 1. The generated SVBRDF parameter maps exhibit strong correlations since they share a common latent representation as the start point for decoding. 2. The decoder module does not compromise generating diversity, as we keep the denoising UNet frozen during the training of the PBR decoder. Additionally, we introduced a highlight-aware decoder for albedo map to further enhance regularization.
@@
  • We propose an image-to-image editing scheme that facilitates material editings tasks such as stylization, inpainting, and seamless texture synthesis.
@@ 2.1. Material estimation
Material estimation approaches aim to acquiring material data from real-world measurement under varing viewpoints and lighting conditions. We specifically focus on recent material estimation methods that utilize lightweight capture setup using consumer cameras. For a more comprehensive overview of general appearance modeling, please refer to surverys (Dong, 2019; Weinmann and Klein, 2015; Guarnera et al., 2016).
Methods have been developed to leverage multiple images or video sequences captured by a handheld camera to estimate appearance properties. Due to the limitations of lightweight setups, most approaches still rely on regularization such as handcrafted heuristics for diffuse/specular separation (Riviere et al., 2016; Palma et al., 2012), linear combinations of basis BRDFs (Hui et al., 2017), and sparsity assumption for incident lighting (Dong et al., 2014). Another class of methods focus on reduce the number of input images by leveraging material priors such as stationary materials  (Aittala et al., 2015, 2016), homogeneous or piece-wise materials  (Xu et al., 2016), and spatially sparse materials  (Zhou et al., 2016).
In recent year, deep learning-based methods have shown significant progress in recoverying SVBRDFs from single image (Li et al., 2017; Deschaintre et al., 2018a; Li et al., 2018; Guo et al., 2021, 2023; Henzler et al., 2021). These methods employ deep convolutional neural network to predict plausible SVBRDFs from in-the-wild input images in a feed-forward manner.  Deschaintre et al. (2019) extended single-image based solution to multiple images by latent space max-pooling. More recent work by Gao et al. (2019) introduced a deep inverse rendering pipeline that enables appearance estimation from an arbitrary number of input images. In procedural material modeling,  Hu et al. (2019); Shi et al. (2020); Hu et al. (2022a) proposed to optimize material parameters with fixed node graphs to match input images.  Hu et al. (2022b) introduced a new pipeline that eliminates the need for predefined node graphs. Most recently,  Sartor and Peers (2023) proposed a diffusion-based model to estimate the material properties from a single photograph.
@@ 2.2. Generative models
Recent advancements in text-to-image generation have been mainly driven by diffusion models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Ramesh et al., 2022). Later advancements (Song et al., 2020, 2021; Liu et al., 2022) have explored efficient sampling strategies to significantly reduce the number of required sampling steps, thereby improving image generation performance.  Rombach et al. (2022) proposed Latent Diffusion Model (LDM) to perform denoising process in learned compact latent space, enabling high-resolution image synthesis and efficient image manipulation.
@@
Integrating multi-modal controllability into text-to-image diffusion model is crucial for creation applications. Recent research (Denis Zavadski and Rother, 2023; Zhang et al., 2023; Ye et al., 2023; Hu et al., 2022c; Mou et al., 2023) has focued on lightweight multi-modal controllability without the requirements of extensive data and high computational power. Hu et al. (2022c) introduces a fine-tuning strategy using low-rank matrices, enabling domain-specific adaptation. Zhang et al. (2023) proposed ControlNet, adding spatial conditioning to diffusion models for precise generation control. Ye et al. (2023) presented a lightweight framework enhancing diffusion models with image prompts using a decoupled cross-attention mechanism.
@@
Guo et al. (2020) proposed an unconditional MaterialGAN for synthesizing SVBRDFs from random noise. The learned latent space facilitates efficient material estimation in inverse renderings. Zhou et al. (2022) developed a StyleGAN2-based model, conditioned by spatial structure and material category, for tileable material synthesis. These GAN-based methods shows advantages in generating high-resolution and visually compelling materials. However, their diversity is constrained by training instability of GANs and the limited range of training datasets. In procedural material generation, Guerrero et al. (2022) first introduced a transformer-based autoregressive model. Later work by Hu et al. (2023) proposed a multi-model node graph generation architecture for creating high-quality procedural materials, guided by both text and image inputs. While procedural representations is compact and resolution-independent, they are limited to stationary patterns and cannot create arbitrary styles.
In concurrent work, Vecchio et al. (2023) introduced ControlMat, a diffusion-based material generative model, capable of generating tileable materials using text and single photograph as input. This model was trained on a synthetic material dataset comprising $126,000$ samples, derived from $8,615$ raw material graphs. While quite large in material domain, this dataset is relatively small compared to the billions of text-images pairs used in text-to-image diffusion model training. This scale discrepancy leads to constrained diversity. Furthermore, this work only supports guidance of text and single photograph, limiting the scenarios range.
In contrast, our method significantly enhance material generation diversity through the efficient integration of pretrained diffusion models with material priors. We also provide a variety of user-friendly controls for guiding the generation process, expanding the scope and flexibility of applications.
@@ 2.3. Text-to-3D Generation
Transitioning 2D text-to-image approaches to 3D generation presents significant challenges, mainly due to lacking large-scale labeled 3D datasets. Recent approaches (Poole et al., 2022; Wang et al., 2023; Lin et al., 2023; Tang et al., 2023) have explored text-to-3D generation without the dependency of 3D data. (Poole et al., 2022) integrates Score Distillation Sampling (SDS) with text-to-image diffusion models. Wang et al. (2023) further improved the quality and diversity by introducing Variational Score Distillation (VSD). The development of large-scale 3D datasets (Deitke et al., 2023) enabled direct learning from 3D data (Liu et al., 2023; Shi et al., 2023). However, current 3D generation methods mainly focus on geometry modeling and fail to producing high-quality, disentangled materials.
Park et al. (2018) introduced a neural method to assign materials from predefined set to different part of a 3D shape. Extending this, Hu et al. (2022d) employs a translation network to establish the correspondence between 2D exemplar image and 3D shape. This allows for extracting material cues from 2D images and selecting optimal materials from a candidate pool using a perceptual metric. However, these methods are constrained by the variety of their predefined material assets and lack the ability to transfer complex spatial patterns from 2D exemplar to 3D shapes. In contrast, our generative model can produce diverse materials and effectively transfer spatial structures from 2D exemplar images to 3D models, showcasing a significant advancement in material assignment.
@@ 3.1. Overview
Figure 2. Overview of DreamPBR: The denoising UNet in our Material LDM is trained with only albedo textures (upper left) and a PBR Decoder with Highlight Aware Decoder is used to transform albedo textures to other physically-based textures (middle right). In the blue box on the left, we presents three individual control modules: Pixel Control, Style Control, and Shape Control, whose results under controls are shown on the lower right. Besides, an addtional Rendering-aware-super-resolution module is given for higher-quality textures (upper right).
@@
The core of our system is encased in a red box, a Material LDM, which is capable of decoding high-quality SVBRDF maps through my PBR decoder and Highlight-aware decoder. Paired with our Rendering-aware super-resolution, high-resolution texture maps can be achieved. In the inference phase, DreamPBR supports multimodal control generation as shown in the left purple box. It can guide the inference process to generate controllable texture maps with input pixel control information and style control information. Additionally, it enables the application of materials to 3D objects through integration with 3D shapes, example images, and existing large language models such as GPT.
@@
The goal of our method is to generate spatially-varying materials which are represented by Cook-Torrance microfacet BRDF model with GGX normal distribution function  (Walter et al., 2007). Specifically, we use metallic-based PBR workflow and represent surface reflectance properties as albedo map $\mathcal{P}$, normal map $\mathcal{N}$, roughness map $\mathcal{R}$, and metallic map $\mathcal{M}$.
@@
The core generative module of our framework is material Latent Diffusion Model (material LDM), which takes textual description $T$ as inputs to encode high-dimensional surface reflectance properties into a compact latent representation $z$. This representation effectively compresses complex material data and guides the SVBRDF decoder in reconstructing detailed SVBRDF maps (i.e. albedo, normal, roughness, and metallic) $S=\{\mathcal{P},\mathcal{N},\mathcal{R},\mathcal{M}\}$. Our critical observation is that while pre-trained text-to-image diffusion models can capture a wide range of natural images that fulfill the diversity needs of material generation, their flexibility often leads to less plausible materials due to the absence of material priors. Instead of training material LDM from scratch with limited material data, we opted to fine-tune a pre-trained text-to-image diffusion model with target material data. This strategy effectively tailors the model from a broad image domain to a specific material domain, ensuring both diversity and authenticity of output.
Our text-to-material framework seamlessly integrates three types of control modules to enhance material generation capabilities. First, we introduce Pixel Control module $G_{P}$ that takes pixel-aligned inputs (e.g. sketches, masks), utilizing the ControlNet architecture (Zhang et al., 2023). It adds conditional controls into diffusion models, providing spatial guidance for material generation. Second, we use Style Control module $G_{I}$ to extract image features from input image prompt, which are then utilized to adapt material LDM via cross attention. Third, we propose Shape control module $G_{S}$ to generate SVBRDF maps automatically for a given segmented 3D shape. This module can leverage large language models to generate text prompts corresponding to different parts of input shape. It also supports taking a 2D photo exemplar as additional input, enabling the generation of material maps for each segmented part, guided by the segmented 2D image. In the rest of current section, we will dive into the key components of our framework. Section 3.2 introduce our core text-to-material module that enables tileable, diverse material generation. Next, Section 3.3 describes the SVBRDF decoder, responsible for reconstructing high-resolution SVBRDF maps from a unifed latent space. Finally, Section 3.4 discusses the Multi-modal control module, providing image and 3D control capabilities to the diffusion model.

### v2

@@ Abstract.
Prior material creation methods had limitations in producing diverse results mainly because reconstruction-based methods relied on real-world measurements and generation-based methods were trained on relatively small material datasets. To address these challenges, we propose DreamPBR, a novel diffusion-based generative framework designed to create spatially-varying appearance properties guided by text and multi-modal controls, providing high controllability and diversity in material generation. The key to achieving diverse and high-quality PBR material generation lies in integrating the capabilities of recent large-scale vision-language models trained on billions of text-image pairs, along with material priors derived from hundreds of PBR material samples. We utilize a novel material Latent Diffusion Model (LDM) to establish the mapping between albedo maps and the corresponding latent space. The latent representation is then decoded into full SVBRDF parameter maps using a rendering-aware PBR decoder. Our method supports tileable generation through convolution with circular padding. Furthermore, we introduce a multi-modal guidance module, which includes pixel-aligned guidance, style image guidance, and 3D shape guidance, to enhance the control capabilities of the material LDM. We demonstrate the effectiveness of DreamPBR in material creation, showcasing its versatility and user-friendliness on a wide range of controllable generation and editing applications.
@@
Figure 1. DreamPBR, an innovative material generation framework, enables personalized creation with multi-modal controls. We present various controls such as text descriptions (4, 5, 6, 7), binary images (2, 9, 10), RGB images (3, 11, 12, 13), segmented geometry (8), and their combination (1) in this figure. The high-quality and tileable textures from DreamPBR show high applicability in different objects.
@@ 1. Introduction
Recent approaches have explored material generation (Guo et al., 2020; Zhou et al., 2022) using Generative Adversarial Networks (GANs) (Goodfellow et al., 2014). However, these methods are typically trained on hundreds to thousands of materials, which pales in comparison to the billions of images used in large-scale Language-Image generative models. The dataset capacity restricted their generating diversity. Furthermore, GAN-based methods also had training challenges including unstable training, mode collapse, and scalability issues with large datasets. On the other hand, diffusion models (Ho et al., 2020; Rombach et al., 2022) have shown significant advancements, exhibiting advantages in scalability and diversity. Recent advances (Poole et al., 2022; Wang et al., 2023) leverage 2D diffusion models before generating 3D content. However, these methods mainly focus on implicit representation or textured mesh, lacking the capability to disentangle physically based material and illumination.
@@
The key idea of our method is to integrate pretrained 2D text-to-image diffusion models (Rombach et al., 2022) with material priors to generate high-fidelity and diverse materials. While 2D text-to-image Latent Diffusion Models (LDM) excel in generating natural images, they had challenges in producing spatially-varying physically-based material maps due to the large domain gaps between natural images and materials. Consequently, adapting pretrained 2D diffusion models into the material domain, while preserving both quality and diversity, is a non-trivial research task. We introduce a novel material LDM which is learned by a two-stage strategy to address this challenge. In the initial stage, we observed albedo map is a specialized RGB image and stores spatially-varying surface reflectance by RGB pixel values. We transfer the pretrained LDM from the text-to-image domain to the text-to-albedo domain using fine-tuning, which can be regarded as the distillation from a large source domain (natural images) to a relatively small target domain (albedo texture maps) by leveraging the target domain priors. In the subsequent stage, we leverage a PBR decoder to reconstruct SVBRDFs from the latent space of albedo maps learned in the former stage. The reasons that we employ a decoder-only architecture for SVBRDFs generation are: 1. The generated SVBRDF parameter maps exhibit strong correlations since they share a common latent representation as the starting point for decoding. 2. The decoder module does not compromise generating diversity, as we keep the denoising UNet frozen during the training of the PBR decoder. Additionally, we introduced a highlight-aware decoder for the albedo map to further enhance regularization.
@@
  • We propose an image-to-image editing scheme that facilitates material editing tasks such as stylization, inpainting, and seamless texture synthesis.
@@ 2.1. Material estimation
Material estimation approaches aim to acquire material data from real-world measurements under varying viewpoints and lighting conditions. We specifically focus on recent material estimation methods that utilize lightweight capture setups using consumer cameras. For a more comprehensive overview of general appearance modeling, please refer to surveys (Dong, 2019; Weinmann and Klein, 2015; Guarnera et al., 2016).
Methods have been developed to leverage multiple images or video sequences captured by a handheld camera to estimate appearance properties. Due to the limitations of lightweight setups, most approaches still rely on regularization such as handcrafted heuristics for diffuse/specular separation (Riviere et al., 2016; Palma et al., 2012), linear combinations of basis BRDFs (Hui et al., 2017), and sparsity assumption for incident lighting (Dong et al., 2014). Another class of methods focuses on reducing the number of input images by leveraging material priors such as stationary materials  (Aittala et al., 2015, 2016), homogeneous or piece-wise materials  (Xu et al., 2016), and spatially sparse materials  (Zhou et al., 2016).
In recent years, deep learning-based methods have shown significant progress in recovering SVBRDFs from single image (Li et al., 2017; Deschaintre et al., 2018a; Li et al., 2018; Guo et al., 2021, 2023; Henzler et al., 2021). These methods employ deep convolutional neural network to predict plausible SVBRDFs from in-the-wild input images in a feed-forward manner.  Deschaintre et al. (2019) extended a single-image-based solution to multiple images by latent space max-pooling. More recent work by Gao et al. (2019) introduced a deep inverse rendering pipeline that enables appearance estimation from an arbitrary number of input images. In procedural material modeling,  Hu et al. (2019); Shi et al. (2020); Hu et al. (2022a) proposed to optimize material parameters with fixed node graphs to match input images.  Hu et al. (2022b) introduced a new pipeline that eliminates the need for predefined node graphs. Most recently,  Sartor and Peers (2023) proposed a diffusion-based model to estimate the material properties from a single photograph.
@@ 2.2. Generative models
Recent advancements in text-to-image generation have been mainly driven by diffusion models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Ramesh et al., 2022). Later advancements (Song et al., 2020, 2021; Liu et al., 2022) have explored efficient sampling strategies to significantly reduce the number of required sampling steps, thereby improving image generation performance.  Rombach et al. (2022) proposed Latent Diffusion Model (LDM) to perform the denoising process in learned compact latent space, enabling high-resolution image synthesis and efficient image manipulation.
@@
Integrating multi-modal controllability into a text-to-image diffusion model is crucial for creation applications. Recent research (Denis Zavadski and Rother, 2023; Zhang et al., 2023; Ye et al., 2023; Hu et al., 2022c; Mou et al., 2023) has focused on lightweight multi-modal controllability without the requirements of extensive data and high computational power. Hu et al. (2022c) introduces a fine-tuning strategy using low-rank matrices, enabling domain-specific adaptation. Zhang et al. (2023) proposed ControlNet, adding spatial conditioning to diffusion models for precise generation control. Ye et al. (2023) presented a lightweight framework enhancing diffusion models with image prompts using a decoupled cross-attention mechanism.
@@
Guo et al. (2020) proposed an unconditional MaterialGAN for synthesizing SVBRDFs from random noise. The learned latent space facilitates efficient material estimation in inverse renderings. Zhou et al. (2022) developed a StyleGAN2-based model, conditioned by spatial structure and material category, for tileable material synthesis. These GAN-based methods show advantages in generating high-resolution and visually compelling materials. However, their diversity is constrained by the training instability of GANs and the limited range of training datasets. In procedural material generation, Guerrero et al. (2022) first introduced a transformer-based autoregressive model. Later work by Hu et al. (2023) proposed a multi-model node graph generation architecture for creating high-quality procedural materials, guided by both text and image inputs. While procedural representations are compact and resolution-independent, they are limited to stationary patterns and cannot create arbitrary styles.
In concurrent work, Vecchio et al. (2023) introduced ControlMat, a diffusion-based material generative model, capable of generating tileable materials using text and a single photograph as input. This model was trained on a synthetic material dataset comprising $126,000$ samples, derived from $8,615$ raw material graphs. While quite large in the material domain, this dataset is relatively small compared to the billions of text-image pairs used in text-to-image diffusion model training. This scale discrepancy leads to constrained diversity. Furthermore, this work only supports guidance of text and single photograph, limiting the scenarios range.
In contrast, our method significantly enhances material generation diversity through the efficient integration of pretrained diffusion models with material priors. We also provide a variety of user-friendly controls for guiding the generation process, expanding the scope and flexibility of applications.
@@ 2.3. Text-to-3D Generation
Transitioning 2D text-to-image approaches to 3D generation presents significant challenges, mainly due to lacking large-scale labeled 3D datasets. Recent approaches (Poole et al., 2022; Wang et al., 2023; Lin et al., 2023; Tang et al., 2023) have explored text-to-3D generation without the dependency of 3D data. (Poole et al., 2022) integrates Score Distillation Sampling (SDS) with text-to-image diffusion models. Wang et al. (2023) further improved the quality and diversity by introducing Variational Score Distillation (VSD). The development of large-scale 3D datasets (Deitke et al., 2023) enabled direct learning from 3D data (Liu et al., 2023; Shi et al., 2023). However, current 3D generation methods mainly focus on geometry modeling and fail to produce high-quality, disentangled materials.
Park et al. (2018) introduced a neural method to assign materials from a predefined set to different parts of a 3D shape. Extending this, Hu et al. (2022d) employs a translation network to establish the correspondence between 2D exemplar image and 3D shape. This allows for extracting material cues from 2D images and selecting optimal materials from a candidate pool using a perceptual metric. However, these methods are constrained by the variety of their predefined material assets and lack the ability to transfer complex spatial patterns from 2D exemplars to 3D shapes. In contrast, our generative model can produce diverse materials and effectively transfer spatial structures from 2D exemplar images to 3D models, showcasing a significant advancement in material assignment.
@@ 3.1. Overview
Figure 2. Overview of DreamPBR: The denoising UNet in our Material LDM is trained with only albedo textures (upper left) and a PBR Decoder with Highlight Aware Decoder is used to transform albedo textures to other physically-based textures (middle right). In the blue box on the left, we present three individual control modules: Pixel Control, Style Control, and Shape Control, whose results under controls are shown on the lower right. Besides, an additional Rendering-aware-super-resolution module is given for higher-quality textures (upper right).
@@
The goal of our method is to generate spatially-varying materials which are represented by the Cook-Torrance microfacet BRDF model with GGX normal distribution function  (Walter et al., 2007). Specifically, we use metallic-based PBR workflow and represent surface reflectance properties as albedo map $\mathcal{P}$, normal map $\mathcal{N}$, roughness map $\mathcal{R}$, and metallic map $\mathcal{M}$.
@@
The core generative module of our framework is the material Latent Diffusion Model (material LDM), which takes textual description $T$ as inputs to encode high-dimensional surface reflectance properties into a compact latent representation $z$. This representation effectively compresses complex material data and guides the SVBRDF decoder in reconstructing detailed SVBRDF maps (i.e. albedo, normal, roughness, and metallic) $S=\{\mathcal{P},\mathcal{N},\mathcal{R},\mathcal{M}\}$. Our critical observation is that while pre-trained text-to-image diffusion models can capture a wide range of natural images that fulfill the diversity needs of material generation, their flexibility often leads to less plausible materials due to the absence of material priors. Instead of training material LDM from scratch with limited material data, we opted to fine-tune a pre-trained text-to-image diffusion model with target material data. This strategy effectively tailors the model from a broad image domain to a specific material domain, ensuring both diversity and authenticity of output.
Our text-to-material framework seamlessly integrates three types of control modules to enhance material generation capabilities. First, we introduce the Pixel Control module $G_{P}$ that takes pixel-aligned inputs (e.g. sketches, masks), utilizing the ControlNet architecture (Zhang et al., 2023). It adds conditional controls into diffusion models, providing spatial guidance for material generation. Second, we use Style Control module $G_{I}$ to extract image features from the input image prompt, which are then utilized to adapt material LDM via cross attention. Third, we propose a Shape control module $G_{S}$ to generate SVBRDF maps automatically for a given segmented 3D shape. This module can leverage large language models to generate text prompts corresponding to different parts of the input shape. It also supports taking a 2D photo exemplar as additional input, enabling the generation of material maps for each segmented part, guided by the segmented 2D image. In the rest of the current section, we will dive into the key components of our framework. Section 3.2 introduce our core text-to-material module that enables tileable, diverse material generation. Next, Section 3.3 describes the SVBRDF decoder, responsible for reconstructing high-resolution SVBRDF maps from a unified latent space. Finally, Section 3.4 discusses the Multi-modal control module, providing image and 3D control capabilities to the diffusion model.